# Overcoming Label Noise for Source-free Unsupervised Video Domain Adaptation

### v1

@@ III-A Problem Definition
Figure 4: Overview of the three stages of our CleanAdapt + TS framework for source-free video domain adaptation, which has three stages. (a) The model ($f_{a}$) is first pre-trained on the labeled source domain videos from $\mathcal{D}_{s}$. For brevity, only the single-stream model is shown here. (b) This source pre-trained model is then used to generate pseudo-labels $\hat{y}$ for the unlabeled target domain videos from $\mathcal{D}_{t}$. Inevitably, these pseudo-labels are noisy due to the domain shift between the source and the target domains. (c) A clean sample selection module is used to select a set $\mathcal{D}_{cl}$ of small-loss samples as potential clean samples. The source pre-trained model is finetuned on these clean samples from $\mathcal{D}_{cl}$ using their corresponding pseudo-labels $\hat{y}$. We repeat this step multiple times. See Sec. IV for implementation details. (Best viewed in color.)
III-B Self-training based Domain Adaptation
@@
In contrast to the adversarial learning based approaches [12, 11, 13], we take the path of self-training [28, 74, 75] primarily due to its simplicity in the adaptation stage. First, we pre-train the 3D CNN models using the labeled source videos from $D_{s}$. Second, we generate pseudo-labels for the unlabeled target dataset $D_{t}$ using the source pre-trained model referred to as pseudo-labels. Third, we retrain the networks $f_{a}$ and $f_{m}$ using the pseudo-labeled target domain videos from $D_{t}$ for adaptation. One of the possibilities is to use all the samples with their corresponding pseudo-labels to retrain the networks. However, pseudo-labels are noisy due to the domain gap between the source domain $D_{s}$ and the target domain $D_{t}$. Retraining $f_{a}$ and $f_{m}$ with all these pseudo-labeled samples from $D_{t}$ leads to a sub-optimal result, as discussed in Section IV. We aim to answer the following question in this paper: How do we choose the pseudo-labeled samples from $D_{t}$ that can help the model in adaptation?
III-C Clean Samples are All You Need
@@
The pseudo-labels contain a large number of samples with correct pseudo-labels (clean samples). For example, there are $\sim$90samples with correct pseudo-labels in the UCF dataset when generated using the HMDB pre-trained networks. Thus, if we can filter out the noisy samples and keep only the clean ones, we can finetune our networks ($f_{a}$ and $f_{m}$) using these clean samples and their corresponding pseudo-labels. Thus, we argue that these clean samples are the ones that can help us in domain adaptation. Now, the important question here is how to separate the clean samples from the noisy ones.
@@
In Figure 3, we observe that deep neural networks learn from clean samples easily and have a hard time learning from noisy samples due to the memorization effect [67]. Thus, samples with low-loss values are the potential clean samples and can be filtered out using the loss as an indicator. In this work, we design two approaches without bells and whistles, CleanAdapt and CleanAdapt + TS, aiming to select the clean samples based on the loss generated by the model against their corresponding pseudo-labels for adaptation. In each epoch of the adaptation stage, we select these clean samples from the target domain and use them to re-train the source pre-trained models $f_{a}$ and $f_{m}$.
@@
There are three key advantages to this: (1) we do not need to modify the overall training regime (e.g., contrastive learning for domain alignment [15, 14, 16]) during adaptation, (2) we do not need to make any domain adaptation specific design choices (e.g., background mixing [16]), and (3) we implicitly design an adaptation method that does not need any source data during the adaptation stage (refer Figure 4).
III-D Source Pre-training
@@
III-E Pseudo-label Generation
@@
III-F CleanAdapt + TS: A Strong Video Adaptation Method
@@
Clean sample selection. To filter out the target domain videos with noisy pseudo-labels, we start with the pseudo-labels generated in Sec. III-E and exploit the relation between the small-loss and the clean samples. We use the source pre-trained models ($f_{a}$, $f_{m}$) to select the clean samples reliably. In each epoch, the videos are first grouped into $|C|$ classes based on their pseudo-labels generated by the model and sorted in ascending order of their cross-entropy loss values computed using the prediction made by the models ($p(x)=\sigma(f_{a}(x)+f_{m}(x))$) and their corresponding pseudo-labels ($\hat{y}$):
@@
where $\sigma(.)$, $D_{cl}$, and $D_{no}$ represent the softmax function, sets of clean samples and noisy samples, respectively. If the pseudo-labels are correct, the model is likely to produce a small loss, and thus, there is a high possibility that the sample belongs to $D_{cl}$. Inspired by [72, 76], we define a hyper-parameter keep-rate $\tau$. For each group, we select $\tau$ proportion of the total number of samples with small losses (See Figure 5). We call this updated dataset of small-loss samples as $D_{cl}\subset D_{t}$ and discard the samples in $D_{no}$. We update the pseudo-labels as follows:
@@
Inspired by the success of the teacher-student framework [29, 30], we adopt it along with our small-loss based clean sample selection to combat the label-noise, and we call this model as CleanAdapt + TS. The teacher and student networks share the same architecture and are initialized with the source pre-trained weights. We use temporally averaged teacher models to select reliable pseudo-labeled target domain videos. To accomplish this, we generate two copies of the models $f_{a}$ and $f_{m}$: one operates as the teacher model ($f_{a}^{T}$, $f_{m}^{T}$), while the other functions as the student model ($f_{a}^{S}$, $f_{m}^{S}$) in their respective modalities. The parameters of the teacher models ($\theta_{a}^{T}$ and $\theta_{m}^{T}$) are not updated through loss back-propagation, rather they are an exponential moving average of student parameters $\theta_{a}^{S}$ and $\theta_{m}^{S}$ respectively. To this end, we create two different versions of the same video $x$ in each modality using two transformations. Let $T_{w}(x)$ and $T_{s}(x)$ denote the weakly and strongly augmented versions of the same video $x\in\mathcal{D}_{t}$ as defined below:
@@
Strong augmentations. To achieve robust augmentation, we implement RandAugment [77] for each video $x$. This method randomly selects three augmentations from a predefined list and applies them to the video $x$.
@@
IV Results and Analysis
IV-A Datasets and Metrics
We consider both first-person and third-person videos for benchmarking our proposed approach. Following [11, 12, 13, 15, 14, 16], we use publicly available UCF101 [78] and HMDB51 [79] for third-person and EPIC-Kitchens [80] for first-person videos. We show experimentally that our approach adapts well to both of these scenarios.
TABLE II: Performance comparison with state-of-the-art video domain adaptation methods on UCF101$\leftrightarrow$ HMDB51. Result for MM-SADA [12] is taken from Kim et al. [15]. The results for our methods are highlighted in gray.
@@ IV-A Datasets and Metrics
UCF normal-$\leftrightarrow$ HMDB. We use the official split released by Chen et al. [11] for UCF $\leftrightarrow$ HMDB to evaluate our CleanAdapt on video domain adaptation. In total, this dataset has 3209 third-person videos with 12 action classes. Specifically, all the videos are a subset of the original UCF101 [78] and HMDB51 [79] datasets with 12 classes common between them. Following [11], we use two settings: UCF101 $\rightarrow$ HMDB51, and HMDB51 $\rightarrow$ UCF101.
UCF normal-$\leftrightarrow$ HMDBsmall. This dataset [81] has 5 shared classes from UCF101 and HMDB51 datasets with a total of 1271 videos.
EPIC-Kitchens. This is the largest video domain adaptation dataset, which contains egocentric videos of fine-grained actions recorded in different kitchens. We follow the official split provided by Munro et al. [12]. This dataset contains videos from the three largest kitchens, i.e., D1, D2, and D3, with 8 common action categories. EPIC-Kitchens has more class imbalance than UCF $\leftrightarrow$ HMDB, making it more challenging [12].
Metrics. We follow the standard protocol defined by [11, 12] to compare our approach with state-of-the-art unsupervised domain adaptation methods [82, 7, 8, 11, 12, 13, 15, 14, 16] in terms of top-1 accuracy. We perform cross-domain retrieval experiments to evaluate the feature space learned by our model before and after adaptation. We report retrieval performance in terms of Recall at $k$ (R@k), implying that if $k$ closest nearest neighbors contain one video of the same class, the retrieval is considered correct.
IV-B Implementation Details
We use the Inception I3D [3] network as the backbone for both RGB and Flow modalities. Following the prior video domain adaptation works [12, 13, 15, 14], we use the Kinetics [83] pre-trained weights to initialize the I3D network. During training, we randomly sample 16 consecutive frames and perform the same data augmentation used in [12, 13, 15] for all our steps. We set the batch size to 48 for both UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets. We pre-compute optical flow using the TV-L1 algorithm [84].
Source pretraining stage. We train the model on the source dataset for 40 and 100 epochs with learning rates $21e-$ 2, and $22e-$ 2 for UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets, respectively. We reduce the learning rate by a factor of 10 after 10, 20 epochs for UCF $\leftrightarrow$ HMDB. For EPIC-Kitchens, we decrease the learning rate by 10 after 50 epochs. We follow [13] for other hyperparameters.
Adaptation stage. We use the source pre-trained weights during the adaptation stage to initialize the I3D [3] network. The network is trained for 100 epochs with learning rates $21e-$ 2 and $32e-$ 3 for UCF $\leftrightarrow$ HMDB and EPIC-Kitchens, respectively. The learning rate is reduced by 10 after 20, 40 epochs for UCF $\leftrightarrow$ HMDB. In the case of EPIC-Kitchens, we reduce the learning rate by 10 after 10, 20 for EPIC-Kitchens. We set the values of momentum parameters $\epsilon$ as 0.99 in all experiments.
Our entire framework is implemented in PyTorch [85] and uses 4 NVIDIA 2080Ti GPUs. On average, training takes around 1 hour for UCF $\leftrightarrow$ HMDB and approximately 7 hours for EPIC-Kitchens datasets.
We now first provide a detailed comparison of our proposed approaches with state-of-the-art video domain adaptation methods on UCF $\leftrightarrow$ HMDB, UCF $\leftrightarrow$ HMDBsmall, and EPIC-Kitchens datasets in Sec. IV-C. We provide some discussions to understand the effect of high-loss samples and over-confident pseudo-labels on the adaptation stage. In Sec. IV-C, we also show and compare the heatmaps generated by our CleanAdapt and source pre-trained model. In Sec. IV-D, we illustrate the impact of the hyperparameter $\tau$ and explore considerations for selecting an appropriate value for it. We also experimentally show the retrieval performance of our proposed approach as well as the source pre-trained model in Sec. IV-F. In Sec. IV-E, we discuss the impact of the teacher-student framework in detail.
IV-C Comparisons to the State-of-the-art Methods
UCF normal-$\leftrightarrow$ HMDB. We present the quantitative results of both of our approaches CleanAdapt and CleanAdapt + TS for UCF $\leftrightarrow$ HMDB dataset in Table II and compare our results with the state-of-the-art unsupervised source-free video domain adaptation approaches [11, 13, 12, 14, 51, 16, 32, 17, 33, 21, 53, 54, 56, 55, 57]. For each approach in Table II, we also report source-only and target-supervised results for fair comparisons wherever applicable. The source-only method refers to the $f_{a}$ and/or $f_{m}$ models trained only on the train split of the source dataset as described in Section III-D and tested directly on the validation split of the target dataset, which serves as a lower bound of the adaptation performance. The target-supervised model is trained and tested on the train and validation split of the target dataset, respectively. This serves as an upper bound to the adaptation performance.
TABLE III: Comparison with state-of-the-art image-based source-free domain adaptation techniques.
Method Backbone UCF normal-$\rightarrow$ HMDB HMDB normal-$\rightarrow$ UCF Source only TRN 72.7 72.2 Kim et al. [51] TRN 69.9 74.9 Li et al. [86] TRN 74.4 67.3 Yang et al. [87] TRN 75.3 76.3 Qiu et al. [88] TRN 75.8 68.2 Source only I3D 80.6 89.3 Yang et al. [23] I3D 86.6 91.4 Liang et al. [22] I3D 82.5 91.9 CleanAdapt I3D 86.1 96.1 CleanAdapt + TS I3D 88.6 96.7

### v2

@@ 3.1 Problem Definition
3.2 Self-training based Domain Adaptation
@@
In contrast to the adversarial learning based approaches [munro2020multi, chen2019temporal, choi2020shuffle], we take the path of self-training primarily due to its simplicity in the adaptation stage. First, we pre-train the 3D CNN models using the labeled source videos from $D_{s}$. Second, we generate pseudo-labels for the unlabeled target dataset $D_{t}$ using the source pre-trained model referred to as pseudo-labels. Third, we retrain the networks $f_{a}$ and $f_{m}$ using the pseudo-labeled target domain videos from $D_{t}$ for adaptation. One of the possibilities is to use all the samples with their corresponding pseudo-labels to retrain the networks. However, pseudo-labels are noisy due to the domain gap between the source domain $D_{s}$ and the target domain $D_{t}$. Retraining $f_{a}$ and $f_{m}$ with all these pseudo-labeled samples from $D_{t}$ leads to a sub-optimal result, as discussed in Section 4. We aim to answer the following question in this paper: How do we choose the pseudo-labeled samples from $D_{t}$ that can help the model in adaptation?
3.3 Clean Samples are All You Need
@@
The pseudo-labels contain a large number of samples with correct pseudo-labels (clean samples). For example, there are $\sim$90% samples with correct pseudo-labels in the UCF dataset when generated using the HMDB pre-trained networks. Thus, if we can filter out the noisy samples and keep only the clean ones, we can finetune our networks ($f_{a}$ and $f_{m}$) using these clean samples and their corresponding pseudo-labels. Thus, we argue that these clean samples are the ones that can help us in domain adaptation. Now, the important question here is how to separate the clean samples from the noisy ones.
@@
In Figure 3, we observe that deep neural networks learn from clean samples easily and have a hard time learning from noisy samples due to the memorization effect [arpit2017closer]. Thus, samples with low-loss values are the potential clean samples and can be filtered out using the loss as an indicator. In this work, we design two approaches without bells and whistles, CleanAdapt and CleanAdapt + TS, aiming to select the clean samples based on the loss generated by the model against their corresponding pseudo-labels for adaptation. In each epoch of the adaptation stage, we select these clean samples from the target domain and use them to re-train the source pre-trained models $f_{a}$ and $f_{m}$.
@@
There are three key advantages to this: (1) we do not need to modify the overall training regime (e.g., contrastive learning for domain alignment [kim2021learning, song2021spatio, sahoo2021contrast]) during adaptation, (2) we do not need to make any domain adaptation specific design choices (e.g., background mixing [sahoo2021contrast]), and (3) we implicitly design an adaptation method that does not need any source data during the adaptation stage (refer Figure 4).
3.4 Source Pre-training
@@
3.5 Pseudo-label Generation
@@
3.6 CleanAdapt + TS: A Strong Video Adaptation Method
@@
Clean sample selection. To filter out the target domain videos with noisy pseudo-labels, we start with the pseudo-labels generated in Sec. 3.5 and exploit the relation between the small-loss and the clean samples. We use the source pre-trained models ($f_{a}$, $f_{m}$) to select the clean samples reliably. In each epoch, the videos are first grouped into $|C|$ classes based on their pseudo-labels generated by the model and sorted in ascending order of their cross-entropy loss values computed using the prediction made by the models ($p(x)=\sigma(f_{a}(x)+f_{m}(x))$) and their corresponding pseudo-labels ($\hat{y}$):
@@
where $\sigma(.)$, $D_{cl}$, and $D_{no}$ represent the softmax function, sets of clean samples and noisy samples, respectively. If the pseudo-labels are correct, the model is likely to produce a small loss, and thus, there is a high possibility that the sample belongs to $D_{cl}$. Inspired by [han2018co], we define a hyper-parameter keep-rate $\tau$. For each group, we select $\tau$ proportion of the total number of samples with small losses (See Figure 5). We call this updated dataset of small-loss samples as $D_{cl}\subset D_{t}$ and discard the samples in $D_{no}$. We update the pseudo-labels as follows:
@@
Inspired by the success of the teacher-student framework [tarvainen2017mean, sohn2020fixmatch], we adopt it along with our small-loss based clean sample selection to combat the label-noise, and we call this model as CleanAdapt + TS. The teacher and student networks share the same architecture and are initialized with the source pre-trained weights. We use temporally averaged teacher models to select reliable pseudo-labeled target domain videos. To accomplish this, we generate two copies of the models $f_{a}$ and $f_{m}$: one operates as the teacher model ($f_{a}^{T}$, $f_{m}^{T}$), while the other functions as the student model ($f_{a}^{S}$, $f_{m}^{S}$) in their respective modalities. The parameters of the teacher models ($\theta_{a}^{T}$ and $\theta_{m}^{T}$) are not updated through loss back-propagation, rather they are an exponential moving average of student parameters $\theta_{a}^{S}$ and $\theta_{m}^{S}$ respectively. To this end, we create two different versions of the same video $x$ in each modality using two transformations. Let $T_{w}(x)$ and $T_{s}(x)$ denote the weakly and strongly augmented versions of the same video $x\in\mathcal{D}_{t}$ as defined below:
@@
Strong augmentations. To achieve robust augmentation, we implement RandAugment [cubuk2020randaugment] for each video $x$. This method randomly selects three augmentations from a predefined list and applies them to the video $x$.
@@
4 Results and Analysis
4.1 Datasets and Metrics
We consider both first-person and third-person videos for benchmarking our proposed approach. Following [song2021spatio, sahoo2021contrast], we use publicly available UCF101 [soomro2012ucf101] and HMDB51 [kuehne2011hmdb] for third-person and EPIC-Kitchens [damen2018scaling] for first-person videos. We show experimentally that our approach adapts well to both of these scenarios.
UCF $\leftrightarrow$ HMDB. We use the official split released by Chen et al[chen2019temporal] for UCF $\leftrightarrow$ HMDB to evaluate our CleanAdapt on video domain adaptation. In total, this dataset has 3209 third-person videos with 12 action classes. Specifically, all the videos are a subset of the original UCF101 [soomro2012ucf101] and HMDB51 [kuehne2011hmdb] datasets with 12 classes common between them. Following [chen2019temporal], we use two settings: UCF101 $\rightarrow$ HMDB51, and HMDB51 $\rightarrow$ UCF101.
UCF $\leftrightarrow$ HMDBsmall. This dataset has 5 shared classes from UCF101 and HMDB51 datasets with a total of 1271 videos.
EPIC-Kitchens. This is the largest video domain adaptation dataset, which contains egocentric videos of fine-grained actions recorded in different kitchens. We follow the official split provided by Munro et al[munro2020multi]. This dataset contains videos from the three largest kitchens, i.e., D1, D2, and D3, with 8 common action categories. EPIC-Kitchens has more class imbalance than UCF $\leftrightarrow$ HMDB, making it more challenging [munro2020multi].
Ego2Exo. We examine and evaluate the effectiveness of our proposed approach for cross-view transfer in videos using the challenging Ego2Exo [kalluri2024lagtran] dataset. This dataset includes videos sourced from the Ego-Exo4D [grauman2024ego] dataset, leveraging their keystep annotations for action labels (e.g., make dough, prepare skillet, etc.). In total, it comprises 4,100 ego videos and 4,986 exo videos for training, while the validation set consists of 3,168 samples from each view.
Metrics. We follow the standard protocol defined by [chen2019temporal, munro2020multi] to compare our approach with state-of-the-art unsupervised domain adaptation methods [kim2021learning, song2021spatio, sahoo2021contrast] in terms of top-1 accuracy. We perform cross-domain retrieval experiments to evaluate the feature space learned by our model before and after adaptation. We report retrieval performance in terms of Recall at $k$ (R@k), implying that if $k$ closest nearest neighbors contain one video of the same class, the retrieval is considered correct.
4.2 Implementation Details
We use the Inception I3D [carreira2017quo] network as the backbone for both RGB and Flow modalities. Following the prior video domain adaptation works [munro2020multi, choi2020shuffle, kim2021learning, song2021spatio], we use the Kinetics [carreira2017quo] pre-trained weights to initialize the I3D network. During training, we randomly sample 16 consecutive frames and perform the same data augmentation used in [munro2020multi, choi2020shuffle, kim2021learning] for all our steps. We set the batch size to 48 for both UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets. We pre-compute optical flow using the TV-L1 algorithm.
Source pretraining stage. We train the model on the source dataset for 40 and 100 epochs with learning rates 1e21e-2, and 2e22e-2 for UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets, respectively. We reduce the learning rate by a factor of 10 after 10, 20 epochs for UCF $\leftrightarrow$ HMDB. For EPIC-Kitchens, we decrease the learning rate by 10 after 50 epochs. We follow [choi2020shuffle] for other hyperparameters.
@@ 4.2 Implementation Details
Method Venue Two-stream? Source-free? Backbone Datasets
UCF $\rightarrow$ HMDB HMDB $\rightarrow$ UCF
Source only [chen2019temporal] I3D 80.6 88.8
TA3N [chen2019temporal] ICCV’19 I3D 81.4 90.5
Target supervised [chen2019temporal] I3D 93.1 97.0
Source only [choi2020shuffle] I3D 80.3 88.8
SAVA [choi2020shuffle] ECCV’20 I3D 82.2 91.2
Target supervised [choi2020shuffle] I3D 95.0 96.8
Source only [munro2020multi] I3D 82.8 90.7
MM-SADA [munro2020multi] CVPR’20 I3D 84.2 91.1
Target supervised [munro2020multi] I3D 98.8 95.0
Source only [song2021spatio] I3D 82.8 89.8
STCDA [song2021spatio] CVPR’21 I3D 83.1 92.1
Target supervised [song2021spatio] I3D 95.8 97.7
Source only [kim2021learning] I3D 82.8 90.7
Kim et al[kim2021learning] ICCV’21 I3D 84.7 92.8
Target supervised [kim2021learning] I3D 98.8 95.0
Source only [sahoo2021contrast] I3D 80.3 88.8
CoMix [sahoo2021contrast] NeurIPS’21 I3D 86.7 93.9
Target supervised [sahoo2021contrast] I3D 95.0 96.8
Source only [xu2022source] TRN 72.8 72.2
ATCoN [xu2022source] ECCV’22 TRN 79.7 $\blacktriangle$ +6.9 85.3 $\blacktriangle$ +13.1
Source only [yin2022mix] I3D 80.8 91.0
MD-DMD [yin2022mix] MM’22 I3D 82.2 92.8
Target supervised [yin2022mix] I3D 98.8 95.0
Source only [huang2022relative] Transformer 81.1 86.8
MTRAN [huang2022relative] MM’22 Transformer 92.2 $\blacktriangle$ +11.1 95.3 $\blacktriangle$ +8.5
Costa et al[da2022dual] WACV’22 I3D 87.8 95.8
Source only ResNet-101 76.4 78.1
@@
Source only [yang2022interact] I3D 85.8 93.5
CIA [yang2022interact] CVPR’22 I3D-TRN 91.9 94.6
Target supervised [yang2022interact] I3D 96.8 99.1
Source only [wei2022unsupervised] I3D 86.1 92.5
TranSVAE [wei2022unsupervised] NeurIPS’23 I3D-TRN 87.8 99.0
Source only I3D 80.6 89.3
CleanAdapt Ours I3D 86.1 $\blacktriangle$ +5.5 96.1 $\blacktriangle$ +6.8
CleanAdapt + TS Ours I3D 88.6 $\blacktriangle$ +8.0 96.7 $\blacktriangle$ +7.4
Target supervised I3D 93.6 98.4
Source only I3D 82.5 91.4
CleanAdapt Ours I3D 89.8 $\blacktriangle$ +7.3 99.2 $\blacktriangle$ +7.8
CleanAdapt + TS Ours I3D 93.6 $\blacktriangle$ +11.1 99.3 $\blacktriangle$ +7.9
Target supervised I3D 95.3 99.3
Adaptation stage. We use the source pre-trained weights during the adaptation stage to initialize the I3D [carreira2017quo] network. The network is trained for 100 epochs with learning rates 1e21e-2 and 2e32e-3 for UCF $\leftrightarrow$ HMDB and EPIC-Kitchens, respectively. The learning rate is reduced by 10 after 20, 40 epochs for UCF $\leftrightarrow$ HMDB. In the case of EPIC-Kitchens, we reduce the learning rate by 10 after 10, 20 for EPIC-Kitchens. We set the values of momentum parameters $\epsilon$ as 0.99 in all experiments.
Our entire framework is implemented in PyTorch and uses 4 NVIDIA 2080Ti GPUs. On average, training takes around 1 hour for UCF $\leftrightarrow$ HMDB and approximately 7 hours for EPIC-Kitchens datasets.
We now first provide a detailed comparison of our proposed approaches with state-of-the-art video domain adaptation methods on UCF $\leftrightarrow$ HMDB, UCF $\leftrightarrow$ HMDBsmall, and EPIC-Kitchens datasets in Sec. 4.3. We provide some discussions to understand the effect of high-loss samples and over-confident pseudo-labels on the adaptation stage. In Sec. 4.3, we also show and compare the heatmaps generated by our CleanAdapt and source pre-trained model. In Sec. 4.4, we illustrate the impact of the hyperparameter $\tau$ and explore considerations for selecting an appropriate value for it. We also experimentally show the retrieval performance of our proposed approach as well as the source pre-trained model in Sec. 4.6. In Sec. 4.5, we discuss the impact of the teacher-student framework in detail.
4.3 Comparisons to the State-of-the-art Methods
UCF $\leftrightarrow$ HMDB. We present the quantitative results of both of our approaches CleanAdapt and CleanAdapt + TS for UCF $\leftrightarrow$ HMDB dataset in Table 2 and compare our results with the state-of-the-art unsupervised source-free video domain adaptation approaches. For each approach in Table 2, we also report source-only and target-supervised results for fair comparisons wherever applicable. The source-only method refers to the $f_{a}$ and/or $f_{m}$ models trained only on the train split of the source dataset as described in Section 3.4 and tested directly on the validation split of the target dataset, which serves as a lower bound of the adaptation performance. The target-supervised model is trained and tested on the train and validation split of the target dataset, respectively. This serves as an upper bound to the adaptation performance.
[…]